# Privately Evaluating Untrusted Black-Box Functions

### v1

@@ Appendix C Small-Diameter Subset Extension Mechanism
5:return $2{\text{\sf P}}_{\tau}[{f}](x)-\frac{3\tau}{2}+Z$ where $Z\sim\text{Lap}{\left({\frac{10}{\varepsilon}}\right)}$
@@
Let $\mathcal{W}$ denote Algorithm 4 and consider a fixed $f:\mathcal{U}^{*}\to[0,r]$, $r\in\mathbb{R}$ and $\varepsilon\in(0,1)$. To prove privacy, it suffices to show that $(2{\text{\sf P}}_{\tau}[{f}](\cdot)-\frac{3\tau}{2})$ has low sensitivity and apply the privacy guarantee of the Laplace mechanism (2.1).
@@
$|(2{\text{\sf P}}_{\tau}[{f}](z)-\frac{3\tau}{2})-(2{\text{\sf P}}_{\tau}[{f}](x)-\frac{3\tau}{2})|=2|{\text{\sf P}}_{\tau}[{f}](z)-{\text{\sf P}}_{\tau}[{f}](x)|\leq 2(4+\frac{3r}{\tau})\leq 10.$
@@
${\text{\sf P}}_{\tau}[{f}]=\mathop{{\mathds{E}}\/}_{\begin{subarray}{c}\ell\sim\{|x|-2\tau,\dots,|x|-\tau\}\\ h\sim{\left\{{|x|-\tau,\dots,|x|}\right\}}\end{subarray}}{\left[{{\text{\sf C}}_{\ell}[{f}](x)}\right]}=\frac{1}{2}(f(x)+|x|-(|x|-\frac{3\tau}{2}))=\frac{1}{2}(f(x)+\frac{3\tau}{2}),$
where $(|x|-\frac{3\tau}{2})$ is the expected value of $\ell$. Hence, $2{\text{\sf P}}_{\tau}[{f}](x)-\frac{3\tau}{2}=f(x)$. The down local guarantee follows from the setting of $\tau=3r$ and by inspection of the definition of ${\text{\sf P}}_{\tau}[{f}]$. ∎
@@ Appendix D Double-Monotonization Privacy Wrapper
[…]

### v2

@@ Appendix C Small-Diameter Subset Extension Mechanism
5:return $2{\text{\sf P}}_{\tau}[{f}](x)-\frac{3(\tau+1)}{2}+Z$ where $Z\sim\text{Lap}{\left({\frac{10}{\varepsilon}}\right)}$
@@
Let $\mathcal{W}$ denote Algorithm 4 and consider a fixed $f:\mathcal{U}^{*}\to[0,r]$, $r\in\mathbb{R}$ and $\varepsilon\in(0,1)$. To prove privacy, it suffices to show that $(2{\text{\sf P}}_{\tau}[{f}](\cdot)-\frac{3(\tau+1)}{2})$ has low sensitivity and apply the privacy guarantee of the Laplace mechanism (2.1).
@@
$|(2{\text{\sf P}}_{\tau}[{f}](z)-\frac{3(\tau+1)}{2})-(2{\text{\sf P}}_{\tau}[{f}](x)-\frac{3(\tau+1)}{2})|=2|{\text{\sf P}}_{\tau}[{f}](z)-{\text{\sf P}}_{\tau}[{f}](x)|\leq 2(4+\frac{3r}{\tau})\leq 10.$
@@
${\text{\sf P}}_{\tau}[{f}]=\mathop{{\mathds{E}}\/}_{\begin{subarray}{c}\ell\sim\{|x|-2\tau,\dots,|x|-\tau\}\\ h\sim{\left\{{|x|-\tau,\dots,|x|}\right\}}\end{subarray}}{\left[{{\text{\sf C}}_{\ell}[{f}](x)}\right]}=\frac{1}{2}(f(x)+|x|-(|x|-\frac{3(\tau+1)}{2}))=\frac{1}{2}(f(x)+\frac{3(\tau+1)}{2}),$
where $(|x|-\frac{3(\tau+1)}{2})$ is the expected value of $\ell$. Hence, $2{\text{\sf P}}_{\tau}[{f}](x)-\frac{3(\tau+1)}{2}=f(x)$. The down local guarantee follows from the setting of $\tau=3r$ and by inspection of the definition of ${\text{\sf P}}_{\tau}[{f}]$. ∎
@@ Appendix D Double-Monotonization Privacy Wrapper
  • If $f$ is Lipschitz, then $\mathcal{W}$ has noise distribution $Z_{\varepsilon}$, for a random variable $Z_{\varepsilon}$ with mean 0 and an exponential tail: for all $k>0$, $\Pr{\big{(}{|Z_{\varepsilon}|>\frac{k}{\varepsilon}}\big{)}}\leq e^{-bk}$.